OpenStudy (anonymous):

Use the given cicumfernce of a circle to find the diameter: 12 inches

OpenStudy (anonymous):

How do u do it?

OpenStudy (anonymous):

\[2\pi r = 12\]

OpenStudy (anonymous):

\[r = 12/2\pi\]

OpenStudy (anonymous):

in my book it gives me the options a.255 meters or b.450 yards

OpenStudy (anonymous):

you will r = 1.91 inches and the diameter is two times the radius which is 3.82 inches. then convert those into meter or yards to satisfy you answer options.